A rental car costs $25 plus a fixed charge per mile driven. The total charge for 320 miles of use was $89. Write an equation for the cost, C(in dollars), in terms of the miles driven ,x.: A rental car costs $25 plus a fixed charge per mile driven. The total charge for 320 miles of use was $89. Write an equation for the cost, C(in dollars), in terms of the miles driven ,x.
Answer by jojo14344(879)

You're eqn will come up:
$Rental charge + ($Rate)(No. of miles driven) = $ Cost
Showing this in figures: x=dollar/mile, C=costin dollars
$25+ ($x/mile)(no. of miles driven) = $Cost
$25 +($x/mile)(320 miles)= $89 ------------------> eqn 1
(x/cross(mile))(320cross(miles))=89-25
x=64/320
x=$0.20
There you go the rental car charge $0.20/mile driven plus $25 flat rate
In doubt? go back eqn 1:
25+(0.20/cross(mile))(320cross(miles))=89
25+64=89
89=89
Thank you,
Jojo